OpenStudy (anonymous):

A and B can do a job in 6 days. if A alone can do the same job in 10 days. How many days B will take to do the same job.

OpenStudy (saifoo.khan):

\[\Large \frac{1}{x}+\frac{1}{10}=\frac{1}{6}\]

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

yes got is

OpenStudy (saifoo.khan):

cool.